# Learning Dense Hand Contact Estimation from Imbalanced Data

### v1

@@ Abstract
Hands are essential to human interaction, and understanding contact between hands and the world can promote comprehensive understanding of their function. Recently, there have been growing number of hand interaction datasets that cover interaction with object, other hand, scene, and body. Despite the significance of the task and increasing high-quality data, how to effectively learn dense hand contact estimation remains largely underexplored. There are two major challenges for learning dense hand contact estimation. First, there exists class imbalance issue from hand contact datasets where majority of samples are not in contact. Second, hand contact datasets contain spatial imbalance issue with most of hand contact exhibited in finger tips, resulting in challenges for generalization towards contacts in other hand regions. To tackle these issues, we present a framework that learns dense HAnd COntact estimation (HACO) from imbalanced data. To resolve the class imbalance issue, we introduce balanced contact sampling, which builds and samples from multiple sampling groups that fairly represent diverse contact statistics for both contact and non-contact samples. Moreover, to address the spatial imbalance issue, we propose vertex-level class-balanced (VCB) loss, which incorporates spatially varying contact distribution by separately reweighting loss contribution of each vertex based on its contact frequency across dataset. As a result, we effectively learn to predict dense hand contact estimation with large-scale hand contact data without suffering from class and spatial imbalance issue. The codes will be released.
@@ 1 Introduction
From infancy, humans rely on physical contact to perceive and interact with the surrounding environment. Among the various mediums of contact, hands play a predominant role, facilitating essential tasks and enabling effective communication. Consequently, developing a robust hand contact estimation model is crucial for advancing our understanding of hand interactions and addressing various challenges grady2021contactopt ; liu2024easyhoi that requires accurate hand contact estimation.
@@
Over the recent years, we have witnessed significant advances in large-scale interaction datasets involving hands. Hand-object interaction datasets hasson2019learning ; chao2021dexycb ; cao2021reconstructing ; hampali2020honnotate ; hampali2022keypoint ; fan2023arctic ; liu2022hoi4d ; kwon2021h2o focused on capturing hand grasps of objects. Hand-hand interaction datasets tzionas2016capturing ; moon2020interhand2 ; moon2024dataset considered two hand interaction from a single person. Hand-face interaction dataset shimada2023decaf investigated on interaction between hand and deformable face. Hand-scene interaction datasets hassan2021populating ; huang2022capturing ; tripathi2023deco extended the scope of interaction to include environments featuring the ground, walls, and large-scale objects. Hand-body interaction dataset yin2023hi4d covered broad range of interaction with human body as part of human-human interaction. Despite the wide range of diversity in the hand interaction datasets with contact annotations hasson2019learning ; chao2021dexycb ; cao2021reconstructing ; hampali2020honnotate ; hampali2022keypoint ; fan2023arctic ; liu2022hoi4d ; kwon2021h2o ; moon2020interhand2 ; tzionas2016capturing ; shimada2023decaf ; hassan2019resolving ; huang2022capturing ; yin2023hi4d , to the best of our knowledge, there have been very few attempts to build hand contact estimation model trained on such diverse hand contact. Based on such diverse sets of hand contact data, we aim to build an effective dense hand contact estimation model.
@@
Figure 1: Two challenges for dense hand contact estimation in the wild. First, hand contact datasets suffer from class imbalance, as the majority of samples contain no contact. Second, spatial imbalance arises because contact points are predominantly concentrated at the fingertips. Due to these issues, models trained on such data struggle to generalize to diverse contact patterns across the hand.
@@
There are two major issues for training a dense hand contact estimation. First, there exists class imbalance between contact and non-contact class. Human participants of the hand interaction datasets are often instructed to perform specific actions such as picking up an object chao2021dexycb ; hampali2020honnotate ; liu2022hoi4d and pointing to specific regions of objects kwon2021h2o or other hands moon2020interhand2 . As most of such actions require precise control with interaction by finger tips, most of the other hand regions (e.g., dorsal area) are overlooked and excluded from being in contact. In Figure 1, we provide comparison of occurrence of contact and non-contact in major hand contact datasets for all vertices of MANO hand model romero2017embodied . This observation shows severe class imbalance issue of hand contact from existing hand contact datasets. Numerically, DexYCB chao2021dexycb dataset has roughly 2.7:1 ratio of non-contact to contact samples. More severe InterHand2.6M moon2020interhand2 dataset has huge imbalance ratio of 19.5:1. Such imbalance is also shown in Decaf shimada2023decaf dataset with 21.7:1 ratio. As stated by numerous studies on data imbalance cao2019learning ; cui2019class ; ridnik2021asymmetric ; lengpolyloss , such data imbalance causes poor performance on underrepresented classes, which is in fact the contacting cases in our scenario. Second, the spatial imbalance issue of hand contact data exists throughout the majority of hand contact datasets. This imbalance arises from the fact that hand interaction predominantly involves the fingertips, which provide a high degree of freedom in both movement and rotation. Due to such high maneuverability and the precise nature of actions from motion capture datasets, most of hand contact from the datasets are significantly skewed to finger tip regions. As shown in Figure 1, the heatmap for hand contact from hand contact datasets are severely skewed towards finger tips. Due to such spatial imbalance issue of hand contact datasets, models directly trained on the datasets struggle to generalize on diverse contact patterns across hand.
@@
Therefore, we propose HACO, a framework that learns dense hand contact estimation from imbalanced data. Initially, HACO is trained on a large-scale, assorted dataset composed of 14 datasets, shown in Table 1, that contains various hand interactions to leverage the power of large-scale training brown2020language ; rombach2022high ; kirillov2023segment ; yang2024depth ; pavlakos2024reconstructing . However, this large-scale dataset still exhibits inherent imbalance in terms of both class and spatial distribution. To address these issues, we introduce two strategies designed to fully exploit the potential of such large-scale data. First, we introduce Balanced Contact Sampling (BCS), which mitigates class imbalance in hand contact datasets by dividing the dataset into multiple sampling groups, each representing different statistics of contact and non-contact instances. To ensure fair representation of all contact statistics, we compute a contact balance score that measures how much each hand contact sample deviates from the dataset-wide average. This metric enables us to assess whether each sample is under-represented or over-represented within the original dataset and to adjust the sampling ratios accordingly, ensuring all contact types are well represented. Second, we present the Vertex-Level Class-Balanced (VCB) loss, which applies spatially varying weights to each vertex to address the spatial imbalance issue in hand contact datasets. Inspired by the core idea of class-balanced (CB) loss cui2019class , which re-weights samples based on class frequency, we adapt this concept to the spatial domain by computing a separate loss weight for each vertex based on how well contact and non-contact cases are represented at that vertex in the dataset. While the original CB loss applies global class-level weighting, our method performs fine-grained, vertex-level weighting, allowing the model to focus on under-represented contact patterns in a spatially aware manner.
@@
  • We introduce HACO, a novel framework that addresses the data imbalance problem in hand contact datasets and enables effective learning of dense hand contact estimation from large-scale data.
@@
  • In the end, HACO achieves substantial performance gains over prior methods across a wide range of hand contact scenarios.
@@ 2 Related works
Dense hand contact estimation. Most of the existing methods that explore in dense hand contact estimation are devoted to dense human-scene contact estimation hassan2021populating ; huang2022capturing ; tripathi2023deco . Given an image of human-scene interaction, such methods estimate contactness for each vertex of 3D human model (i.e., SMPL loper2015smpl , SMPL-X pavlakos2019expressive ). POSA hassan2021populating is based upon the idea that the human-scene contact is largely influenced by the 3D pose of human. Hence, the work employs a conditional variational autoencoder framework sohn2015learning to learn human-scene contact with the 3D vertex position of posed human as condition. BSTRO huang2022capturing takes another approach for estimating human-scene contact, which directly relies on visual input with Transformer-based architecture devlin2019bert . To train the model, they proposed RICH dataset huang2022capturing that annotated dense human-scene contact by capturing 3D geometry of body and scene with multi-view laser scanner in outdoor environment. Lastly, to improve generalization ability of human-scene contact estimation, DECO tripathi2023deco focused on how to annotate human-scene contact in the wild. They introduced simple yet effective contact annotation method that simply asks crowd-sourced annotators to paint 3D human model (i.e., SMPL loper2015smpl ) in neutral T-pose, which significantly enhanced the potential of scaling contact data. Despite numerous attempts on dense human-scene contact estimation, there lacks in-depth investigation on dense hand contact estimation. Furthermore, we found that dense hand contact estimation has unique challenges compared to human-scene contact estimation. First, due to class imbalance of hand contact datasets, the distribution of the datasets are often skewed towards non-contact class, making it non-trivial to train hand contact estimation from the datasets. Second, there exists spatial imbalance issue of hand contact datasets where majority of hand contact occurs in finger tips with insufficient data for hand contact in rest of the regions (e.g., dorsum). We introduce HACO, a dense hand contact estimation method that effectively tackles such data imbalance issues of hand contact datasets.
@@
InterHand2.6M moon2020interhand2 Hand Real / Indoor / Studio 2.6M 27 -
HIC tzionas2016capturing Hand Real / Indoor 36K 1 -
PROX hassan2019resolving Scene Real / Indoor / Studio 100K 20 -
RICH huang2022capturing Scene Real / Outdoor 540K 22 -
Decaf shimada2023decaf Face (Whole-Body) Real / Indoor 100K 8 -
Hi4D yin2023hi4d Whole-Body Real / Indoor / Studio 11K 40 -
3D hand and object interaction reasoning. Dense hand contact reveals key spatial and semantic relationship between 3D hand and object. Initially, dense hand contact was studied in the field of grasping with specific tasks such as grasp contact prediction brahmbhatt2019contactdb , contact-guided hand and object pose refinement grady2021contactopt , and contact-conditioned grasp generation liu2023contactgen . Such works made tremendous impact on robotics, especially how robots can naturally grasp an object. However, their dense hand contact were mostly predicted from 3D geometry of hand and object without exploring how to directly predict the dense hand contact from visual input. Recently, the field of 3D hand and object reconstruction showed potential on the effectiveness of contact for improving both reconstruction accuracy and natural interaction between hand and object. CHOI hu2024learning proposed contact conditioned implicit reconstruction model for 3D hand-held object that utilizes contact as a major cue for predicting 3D object surface geometry. EasyHOI liu2024easyhoi introduced a prior-guided hand-object interaction optimizer that refines initial 3D hand and object using heuristically determined contact regions to improve interaction quality. Our HACO is applicable to these downstream tasks as long as visual input is available. To demonstrate HACO’s effectiveness, we show that our predicted dense contact improves 3D grasp optimization over geometry-driven contact estimation from ContactOpt grady2021contactopt , and 3D reconstruction over previously used heuristic contact regions in EasyHOI liu2024easyhoi .
@@ 3 Method
Figure 2: Overall pipeline of HACO. Our method encodes input image as image tokens with a ViT backbone after patch embedding layers. Given the image tokens along with positional embeddings and a contact token, multiple layers of self-attention Transformer and cross-attention Transformer produce an output token. Lastly, the output token is further processed with a linear layer and added with contact initialization that passes sigmoid layer to output final hand contact prediction.
@@ 3.2 Balanced contact sampling
where higher scores correspond to hand instances $\mathbf{H}_{i}$ whose contact patterns deviate more from the dataset-wide average, enabling balanced grouping across diverse contact statistics that span both contact and non-contact samples. Based on the distribution of contact balance scores $\{s_{i}\}_{i=1}^{N}$, we partition the dataset into $K$ bins $\mathcal{B}_{1},\dots,\mathcal{B}_{K}$ by applying non-linear binning. We first compute the minimum and maximum contact balance scores as $s_{\min}=\min_{i}s_{i}$ and $s_{\max}=\max_{i}s_{i}$, and define $K{+}1$ bin edges using a logarithmic spacing scheme controlled by a curvature parameter $\beta>0$ (e.g., $\beta=5$):
@@
3.3 Vertex-level class balanced loss
As illustrated in Figure 1, hand contact datasets exhibit severe spatial imbalance, with contact heavily concentrated at the fingertips. To address this issue, we propose a vertex-level class-balanced (VCB) loss, which builds upon the class-balanced sigmoid binary cross-entropy loss cui2019class . We first formulate hand contact estimation as a binary classification problem. A simple binary cross-entropy (BCE) loss is defined as $\ell_{\text{BCE}}(y,p)=-y\log(p)-(1-y)\log(1-p)$, where $y\in\{0,1\}$ is the ground-truth label and $p\in[0,1]$ is the predicted probability. When applied over all hand vertices $V$, the overall BCE loss becomes:
@@ 3.3 Vertex-level class balanced loss
This spatial adaptivity enables the VCB loss to effectively mitigate the spatial imbalance inherent in hand contact datasets.

### v2

@@ Abstract
Hands are essential to human interaction, and exploring contact between hands and the world can promote comprehensive understanding of their function. Recently, there have been growing number of hand interaction datasets that cover interaction with object, other hand, scene, and body. Despite the significance of the task and increasing high-quality data, how to effectively learn dense hand contact estimation remains largely underexplored. There are two major challenges for learning dense hand contact estimation. First, there exists class imbalance issue from hand contact datasets where majority of regions are not in contact. Second, hand contact datasets contain spatial imbalance issue with most of hand contact exhibited in finger tips, resulting in challenges for generalization towards contacts in other hand regions. To tackle these issues, we present a framework that learns dense HAnd COntact estimation (HACO) from imbalanced data. To resolve the class imbalance issue, we introduce balanced contact sampling, which builds and samples from multiple sampling groups that fairly represent diverse contact statistics for both contact and non-contact vertices. Moreover, to address the spatial imbalance issue, we propose vertex-level class-balanced (VCB) loss, which incorporates spatially varying contact distribution by separately reweighting loss contribution of each vertex based on its contact frequency across dataset. As a result, we effectively learn to predict dense hand contact estimation with large-scale hand contact data without suffering from class and spatial imbalance issue. The codes are available at https://github.com/dqj5182/HACO_RELEASE.
@@ 1 Introduction
From infancy, humans rely on physical contact to perceive and interact with the surrounding environment. Among the various mediums of contact, hands play a predominant role, facilitating essential tasks and enabling effective communication. Consequently, developing a robust hand contact estimation model is crucial for advancing our understanding of hand interactions and addressing various challenges grady2021contactopt ; liu2024easyhoi that require accurate hand contact estimation.
@@
Over the recent years, we have witnessed significant advances in large-scale interaction datasets involving hands. Hand-object interaction datasets hasson2019learning ; chao2021dexycb ; cao2021reconstructing ; hampali2020honnotate ; hampali2022keypoint ; fan2023arctic ; liu2022hoi4d ; kwon2021h2o focused on capturing hand grasps of objects. Hand-hand interaction datasets tzionas2016capturing ; moon2020interhand2 considered two hand interaction from a single person. Hand-face interaction dataset shimada2023decaf investigated on interaction between hand and deformable face. Hand-scene interaction datasets hassan2021populating ; huang2022capturing ; tripathi2023deco extended the scope of interaction to include environments featuring the ground, walls, and large-scale objects. Hand-body interaction dataset yin2023hi4d covered broad range of interaction with human body as part of human-human interaction. Despite the wide range of diversity in the hand interaction datasets with contact annotations hasson2019learning ; chao2021dexycb ; cao2021reconstructing ; hampali2020honnotate ; hampali2022keypoint ; fan2023arctic ; liu2022hoi4d ; kwon2021h2o ; moon2020interhand2 ; tzionas2016capturing ; shimada2023decaf ; hassan2019resolving ; huang2022capturing ; yin2023hi4d ; xie2024rhobin , to the best of our knowledge, there has been no attempt to build hand contact estimation model trained on such diverse hand contact. Based on such diverse sets of hand contact data, we aim to build an effective dense hand contact estimation model.
[…]
@@
Figure 1: Two challenges for dense hand contact estimation in the wild. First, hand contact datasets suffer from class imbalance, as the majority of vertices contain no contact. Second, spatial imbalance arises because contact points are predominantly concentrated at the fingertips. Due to these issues, models trained on such data struggle to generalize to diverse contact patterns across the hand.
@@
There are two major issues for training a dense hand contact estimation. First, there exists class imbalance between contact and non-contact class. Human participants of the hand interaction datasets are often instructed to perform specific actions such as picking up an object chao2021dexycb ; hampali2020honnotate ; liu2022hoi4d and pointing to specific regions of objects kwon2021h2o or other hands moon2020interhand2 . As most of such actions require precise control with interaction by a small region on hand surface, most of the hand regions are overlooked and excluded from being in contact. In Figure 1, we provide comparison on the occurrence of contact and non-contact in major hand contact datasets for all vertices of MANO hand model romero2017embodied . This observation shows severe class imbalance issue of hand contact from existing hand contact datasets. Numerically, DexYCB chao2021dexycb dataset has roughly 2.7:1 ratio of non-contact to contact vertices. More severe InterHand2.6M moon2020interhand2 dataset has huge imbalance ratio of 19.5:1. Such imbalance is also shown in Decaf shimada2023decaf dataset with 21.7:1 ratio. As stated by numerous studies on data imbalance cao2019learning ; cui2019class ; ridnik2021asymmetric ; lengpolyloss , such data imbalance causes poor performance on underrepresented classes, which is in fact the contacting cases in our scenario. Second, the spatial imbalance issue of hand contact data exists throughout the majority of hand contact datasets. This imbalance arises from the fact that hand interaction predominantly involves the fingertips, which provide a high degree of freedom in both movement and rotation. Due to such high maneuverability and the precise nature of actions from motion capture datasets, most of hand contact from the datasets are significantly skewed to finger tip regions. As shown in Figure 1, the heatmap for hand contact from hand contact datasets are highly concentrated at the finger tips. Due to such spatial imbalance issue of hand contact datasets, models directly trained on the datasets struggle to generalize on diverse contact patterns across hand.
@@
Therefore, we propose HACO, a framework that learns dense hand contact estimation from imbalanced data. Initially, HACO is trained on a large-scale, assorted dataset composed of 14 datasets, shown in Table 1, that contains various hand interactions to leverage the power of large-scale training brown2020language ; rombach2022high ; kirillov2023segment ; yang2024depth ; pavlakos2024reconstructing . However, this large-scale dataset still exhibits inherent imbalance in terms of both class and spatial distribution. To address these issues, we introduce two strategies designed to fully exploit the potential of such large-scale data. First, we introduce Balanced Contact Sampling (BCS), which mitigates class imbalance in hand contact datasets by dividing the dataset into multiple sampling groups, each representing different statistics of contact and non-contact. To ensure fair representation of the contact statistics, we compute a contact balance score that measures how much each hand contact instance deviates from the dataset-wide average. This metric enables us to assess whether each instance is under-represented or over-represented within the original dataset and to adjust the sampling ratios accordingly, ensuring all contact types are well represented. Second, we present the Vertex-Level Class-Balanced (VCB) loss, which applies spatially varying weights to each vertex to address the spatial imbalance issue in hand contact datasets. Inspired by the core idea of class-balanced (CB) loss cui2019class , which re-weights the loss based on class frequency, we adapt this concept to the spatial domain by computing a separate loss weight for each vertex based on how well contact and non-contact are represented at the vertex within the dataset. While the original CB loss applies global class-level weighting, our method performs fine-grained, vertex-level weighting, allowing the model to focus on under-represented contact patterns in a spatially-aware manner.
@@
  • We introduce HACO, a novel framework that addresses the data imbalance problems in hand contact datasets and enables effective learning of dense hand contact estimation from large-scale data.
@@
  • In the end, HACO, the first method trained for dense hand contact estimation on a large-scale dataset, achieves substantial performance gains across diverse contact scenarios.
@@ 2 Related works
Dense hand contact estimation. Most of the existing methods that explore dense hand contact estimation are devoted to dense human-scene contact estimation hassan2021populating ; huang2022capturing ; tripathi2023deco ; nam2024joint . Given an image of human-scene interaction, such methods estimate contactness for each vertex of 3D human model (e.g., SMPL loper2015smpl , SMPL-X pavlakos2019expressive ). POSA hassan2021populating is based upon the idea that the human-scene contact is largely influenced by the 3D pose of human. Hence, the work employs a conditional variational autoencoder framework sohn2015learning to learn human-scene contact with the 3D vertex position of posed human as condition. BSTRO huang2022capturing takes another approach for estimating human-scene contact, which directly relies on visual input with Transformer-based architecture devlin2019bert . To train the model, they proposed RICH dataset huang2022capturing that annotates dense human-scene contact by capturing 3D geometry of body and scene with multi-view laser scanner in both indoor and outdoor environment. Lastly, to improve generalization ability of human-scene contact estimation, DECO tripathi2023deco focused on how to annotate human-scene contact in the wild. They introduced simple yet effective contact annotation method that simply asks crowd-sourced annotators to paint 3D human model (i.e., SMPL loper2015smpl ) in neutral T-pose, which significantly enhanced the potential of scaling contact data. Despite numerous attempts on dense human-scene contact estimation, there lacks in-depth investigation on dense hand contact estimation. Furthermore, we found that dense hand contact estimation has unique challenges compared to human-scene contact estimation. First, due to class imbalance of hand contact datasets, the distribution of the datasets are often skewed towards non-contact class, making it non-trivial to train hand contact estimation from the datasets. Second, there exists spatial imbalance issue of hand contact datasets where majority of hand contact occurs in finger tips with insufficient data for hand contact in rest of the regions (e.g., dorsum). We introduce HACO, a dense hand contact estimation method that effectively tackles such data imbalance issues of hand contact datasets.
@@
InterHand2.6M moon2020interhand2 Hand Real / Indoor / Studio 2.6M 27
HIC tzionas2016capturing Hand Real / Indoor 36K 1
PROX hassan2019resolving Scene Real / Indoor / Studio 100K 20
RICH huang2022capturing Scene Real / Outdoor 540K 22
Decaf shimada2023decaf Face (Whole-Body) Real / Indoor 100K 8
Hi4D yin2023hi4d Whole-Body Real / Indoor / Studio 11K 40
3D hand and object interaction reasoning. Dense hand contact reveals key spatial and semantic relationship between 3D hand and object. Initially, dense hand contact was studied in the field of grasping with specific tasks such as grasp contact prediction brahmbhatt2019contactdb , contact-guided hand and object pose refinement grady2021contactopt , and contact-conditioned grasp generation liu2023contactgen . Such works made tremendous impact on robotics, especially on how robots can naturally grasp an object. However, their dense hand contact were mostly predicted from 3D geometry of hand and object without exploring how to directly predict the dense hand contact from visual input. Recently, the field of 3D hand and object reconstruction showed potential on the effectiveness of contact for improving both reconstruction accuracy and natural interaction between hand and object. ContactOpt grady2021contactopt proposed a differentiable contact optimization that optimizes hand and object poses inferred from image-based methods with estimated contact to achieve natural interaction. EasyHOI liu2024easyhoi introduced a prior-guided hand-object interaction optimizer that refines initial 3D hand and object using heuristically determined contact regions to improve interaction quality. Our HACO is applicable to these downstream tasks as long as visual input is available. To demonstrate HACO’s effectiveness, we show that our predicted dense contact improves 3D grasp optimization over geometry-driven contact estimation from ContactOpt grady2021contactopt , and 3D reconstruction over the previously used heuristic contact regions in EasyHOI liu2024easyhoi .
@@ 3 Method
Figure 2: Overall pipeline of HACO. Our method encodes input image as image tokens with a ViT backbone after patch embedding layers. Given the image tokens along with positional embeddings and a contact token, multiple layers of self-attention Transformer and cross-attention Transformer produce an output token. Lastly, the output token is passed through a linear layer and combined with the contact initialization, followed by a sigmoid activation to output the final hand contact prediction.
@@ 3.2 Balanced contact sampling
where higher scores correspond to hand instances $\mathbf{H}_{i}$ whose contact patterns deviate more from the dataset-wide average, enabling balanced grouping across diverse contact statistics that span both contact and non-contact samples. Based on the distribution of contact balance scores $\{s_{i}\}_{i=1}^{N}$, we partition the dataset into $K$ bins $\mathcal{B}_{1},\dots,\mathcal{B}_{K}$ by applying non-linear binning. We first compute the minimum and maximum contact balance scores as $s_{\min}=\min_{i}s_{i}$ and $s_{\max}=\max_{i}s_{i}$, and define K+1K{+}1 bin edges using a logarithmic spacing scheme controlled by a curvature parameter $\beta>0$ (e.g., $\beta=5$):
@@
3.3 Vertex-level class-balanced loss
As illustrated in Figure 1, hand contact datasets exhibit severe spatial imbalance, with contact heavily concentrated at the fingertips. To address this issue, we propose a vertex-level class-balanced (VCB) loss, which builds upon the class-balanced (CB) sigmoid binary cross-entropy loss cui2019class . We first formulate hand contact estimation as a binary classification problem. A simple binary cross-entropy (BCE) loss is defined as $\ell_{\text{BCE}}(y,p)=-y\log(p)-(1-y)\log(1-p)$, where $y\in\{0,1\}$ is the ground-truth label and $p\in[0,1]$ is the predicted probability. When applied over all hand vertices $V$, the overall BCE loss becomes:
@@ 3.3 Vertex-level class-balanced loss
This spatial adaptivity enables the VCB loss to effectively mitigate the spatial imbalance inherent in hand contact datasets. The VCB loss is applied through a progressive weighting strategy that combines the CB loss and VCB loss throughout training. At the beginning of training, only the CB loss is applied, while the weight of the VCB component increases linearly with each epoch. The VCB weight reaches a fixed maximum at the final epoch, resulting in a smooth transition from global to vertex-adaptive supervision. Lastly, we include a lightweight false-positive regularization term to suppress overly confident contact predictions for non-contact vertices.